# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ 5.3 Return amplitudes and probabilities
A second application of the evolution operator (5.6) is to yield the solution $U(t)|\psi(0)\rangle$ of the Schrödinger equation (4.3) for any initial condition $|\psi(0)\rangle$. As in Sec. 4, we choose $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$ to be an eigenstate of the initial inhomogeneous Hamiltonian, and investigate the return amplitude (4.1). This can again be done in terms of $2\times 2$ matrices, and the return amplitude can in fact be evaluated for any time $t$ (as opposed to $t=T$ at the end of one cycle). In terms of the convenient parameter $h_{k}\equiv h+\tfrac{c}{24}(k^{2}-1)$, one finds (see Appendix D)
$\langle\psi(0)|\psi(t)\rangle=e^{i\pi ck^{2}t/12L}u_{t}^{-2h_{k}/k}$ (5.13)
@@
$u_{t}$ $=\cos(k\Omega t/2)\cos(k\omega t/2)-2\pi\frac{\cosh(2\lambda)-L/T}{L\Omega}\sin(k\Omega t/2)\sin(k\omega t/2)$
$\quad+i\cosh(2\lambda)\biggl[\cos(k\Omega t/2)\sin(k\omega t/2)+2\pi\frac{\operatorname{sech}(2\lambda)-L/T}{L\Omega}\sin(k\Omega t/2)\cos(k\omega t/2)\biggr]$ (5.14)
@@
expressed in terms of the frequency $\Omega$ in Eq. (5.8). The overall phase of $\langle\psi(0)|\psi(t)\rangle$ can be extracted by picking the branch of $\arg(u_{t})$ consistent with the limiting case $\lambda=0$, for which the phase is purely dynamical. This generalizes previous results for the return probability in $\mathfrak{sl}(2,\mathbb{R})$-driven CFT [68], which were restricted to piecewise-constant time-dependent Hamiltonians. Furthermore, the return probability was recently studied for continuous Floquet drives in [69], which relied on Floquet perturbation theory, while the result (5.13) is exact and uses no perturbation theory.
@@
As before, one is most interested in the return amplitude (4.1) after one period, in the adiabatic limit $L/T\to 0$. Taking $t=T$ in Eq. (5.13) and expanding (5.14) in that limit yields
$\langle\psi(0)|\psi(T)\rangle=\exp\Bigl[2\pi i\Bigl(-h_{0}\frac{T}{L}+h_{k}\bigl[\cosh(2\lambda)-1\bigr]+h_{k}\frac{\sinh^{2}(2\lambda)}{2}\frac{L}{T}\Bigr)\Bigr]\Bigl[1+O\bigl(L^{2}/T^{2}\bigr)\Bigr]$ (5.15)
@@
where the exponent manifestly involves a dynamical phase and a Berry phase, plus a first nonadiabatic correction. The Berry phase, in particular, is consistent with Eq. (4.2) for a rotating drive in $\mathrm{SL}(2,\mathbb{R})$ [53].
@@
Note that the return amplitude (5.15) has unit norm up to $L^{2}/T^{2}$ corrections, so it is those corrections that are responsible for the nontrivial return probability $|\langle\psi(0)|\psi(T)\rangle|^{2}$. As in Sec. 4, the corrections are most conveniently found by expanding in powers of the adiabatic parameter (4.17) rather than $L/T$. One finds
$|\langle\psi(0)|\psi(T)\rangle|^{2}=1-\delta^{2}\frac{2h_{k}}{k}\sinh^{2}(2\lambda)\sin^{2}(\pi k/\delta)+O\bigl(\delta^{3}\bigr),$ (5.16)
@@
which shows that there is an oscillating tail in the return probability beyond the adiabatic limit, as in Fig. 6. This result is a special case of the general adiabatic expansion (4.21). A single Fourier mode $k$ now appears in the quantum metric (4.22), since in fact the vector field $\mathbf{Y}_{t}=Y_{t}(\varphi)\partial_{\varphi}$ of Eq. (4.20) is now explicitly given by $Y_{t}(\varphi)=\sinh(2\lambda)\bigl[\sin(k\varphi)-\sin(k[\varphi+\Omega t])\bigr]$. The oscillations of Eq. (5.16) with $T$ can then be understood by referring to Fig. 5: the micromotion around the leading adiabatic approximation has a (high) frequency $\Omega$ of its own, which may or may not resonate with the (low) frequency $\omega$ of the drive. At resonance, when $\Omega T\in 2\pi\mathbb{Z}$, the initial and final rays coincide.
@@
In Fig. 6, we also plot the corresponding result for a periodic free-fermion chain at half filling with $N$ sites and driven, inhomogeneous nearest-neighbor couplings of the form
$J_{j}(t)=\cosh(2\lambda)-\sinh(2\lambda)\cos(k(2\pi j/N-\omega t)).$ (5.17)
@@
The return probability $|\langle\psi_{N}(0)|\psi_{N}(T)\rangle|^{2}$ of the lattice quantum state is computed numerically by approximating the evolution operator with a finite number of steps and using exact diagonalization for each time step; see Appendix E. The result exhibits remarkable agreement with the analytical formula that can be deduced from the square of Eq. (5.13). This is true even far away from the adiabatic regime, as long as the strength of the perturbation (controlled by $\lambda$) is sufficiently weak. Importantly, CFT results for the return probability are even applicable to lattice systems with a relatively small number of lattice sites, as shown in Fig. 6(b). This suggests that our dynamical probes for quantum geometry could potentially be implemented in quantum simulators with fewer than a hundred sites.
@@
Figure 6: (a) Return probability after time evolving under the Hamiltonian (5.3) for one full cycle. The results are plotted as a function of $T/L$ for $\lambda=0.15,0.2,0.25$ (green to purple). [The parameters are $L=200$, $k=2$, $c=1$, $h=0$ and periodic boundary conditions are imposed.] We compare the exact CFT result, given by the square of Eq. (5.13), with numerical lattice calculations for a free-fermionic chain of length $N=L$ (unit lattice spacing) at half filling for inhomogeneous couplings of the form (5.17). (b) Finite-size scaling of the numerical lattice calculations close to the adiabatic limit, and comparison with the exact CFT result, for $k=2$, $\lambda=0.3$, and different system sizes $L$ ranging from 60 to 180 sites. As in Fig. 1, the agreement between numerics and analytical results is striking.
6 Discussion
@@
This work explored the infinite-dimensional quantum geometry of 1+1D gapless many-body quantum systems through the lens of their low-energy dynamics, modeled as driven conformal field theories (CFTs). For slowly driven systems, we showed that the return probability probes the quantum metric through oscillations as a function of the driving frequency: see Figs. 1 and 6. In the complementary limit of arbitrarily fast but weak drives, absorption rates and linear-response coefficients respectively give access to the metric and the Berry curvature of the same quantum geometry. These predictions are universal in the sense that they apply to any 1+1D gapless quantum system with an emergent conformal symmetry at low energies. They hold for both free and strongly interacting systems, and they only depend on the central charge of the effective theory. This is supported by a remarkable agreement between our field-theoretic results and numerics for gapless lattice models, making the observation of the Virasoro quantum geometry plausible in state-of-the-art quantum simulation experiments.
Some future research directions are worth mentioning:
@@
Geometric formulation of Floquet drives. It was recently understood that Floquet theory can be formulated from a quantum geometric perspective [70], by decomposing the Floquet evolution operator into a purely dynamical part (the ‘average-energy operator’) and a geometric part. It would be interesting to compute the average-energy operator for our periodic drives, and understand the link between the Virasoro Berry phases (4.2) [53] and those of [70].
@@
Higher dimensions. Our study of the quantum geometry of 1+1D gapless many-body quantum systems relied on the Virasoro algebra, allowing us to probe an infinite-dimensional parameter space (2.12). While this algebraic structure is absent in higher-dimensional CFTs, one may still design nonequilibrium probes of the quantum geometry associated with the global conformal group SO(d+1,2)(d+1,2) of gapless many-body systems in dd+1 spacetime dimensions. Indeed, given that the approach of this work holds for any continuous symmetry group, it could naturally be applied to higher-dimensional CFTs.
@@
Quantum geometry of mixed states. Our study focused on probes of the quantum geometry of pure states, namely the Virasoro coherent states (2.17). The relevant quantum metric then is the Fubini-Study metric of Sec. 2.3. A natural next step is to consider mixed states instead, and investigate their quantum geometry. In the strict adiabatic regime, one may expect a suitable generalization of our protocol to probe the Uhlmann phase [71, 72], as a mixed state counterpart of the Berry phase. Including subleading corrections in the adiabatic regime should probe the so-called Bogoliubov-Kubo-Mori metric, which was recently studied for thermal initial states in driven CFTs [33].
Acknowledgments
@@
We thank Mathieu Beauvillain, Marin Bukov, Adolfo del Campo, Raphaël Ducatez, Benoit Estienne, Gian Michele Graf, Gregor Jotzu, Bruno Mera, Shinsei Ryu, and Erik Tonni for inspiring discussions. B.L. acknowledges financial support from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 214461). P.M. acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002).
@@ Appendix A Brief dictionary of CFT formulas
This appendix completes the lightning review of CFT and Virasoro symmetry in Sec. 2. Table 1 summarizes useful CFT formulas in angle coordinates $\varphi$, $\theta$ vs. length coordinates $x\equiv\frac{L}{2\pi}\varphi$, $y\equiv\frac{L}{2\pi}\theta$. Quantities are distinguished by their arguments, $\delta(\varphi)$ and $\delta(x)$ respectively denote the $2\pi$- and $L$-periodic delta functions, a prime denotes differentiation with respect to $\varphi$, and $\oint\textrm{d}\phi$ is shorthand for $\int_{0}^{2\pi}\textrm{d}\phi$. Moreover, $\operatorname{Diff}(S^{1})$ denotes the universal cover of the group of orientation-preserving diffeomorphisms of the circle $S^{1}$. $\operatorname{Vect}(S^{1})$ denotes the corresponding Lie algebra of real, smooth vector fields on $S^{1}$, and $\operatorname{Vect}(S^{1})\otimes\mathbb{C}$ denotes its complexification.
@@
Virasoro algebra ${\mathcal{T}}(\varphi)=\frac{1}{2\pi}\underset{n\in\mathbb{Z}}{\sum}e^{in\varphi}\bigl(L_{n}-\frac{c}{24}\delta_{n,0}\bigr)$ ${\mathcal{T}}(x)=\bigl(\frac{2\pi}{L}\bigr)^{2}{\mathcal{T}}(\varphi)=\frac{2\pi}{L^{2}}\underset{n\in\mathbb{Z}}{\sum}e^{i\frac{2\pi}{L}nx}\bigl(L_{n}-\frac{c}{24}\delta_{n,0}\bigr)$
$L_{m}=\oint\textrm{d}\varphi\,e^{-im\varphi}\,{\mathcal{T}}(\varphi)+\frac{c}{24}\delta_{m,0}$ satisfying Eq. (2.9)
@@
$\bigl[{\mathcal{T}}(\varphi),{\mathcal{T}}(\theta)\bigr]=-2i\delta^{\prime}(\varphi-\theta){\mathcal{T}}(\theta)$ $\bigl[{\mathcal{T}}(x),{\mathcal{T}}(y)\bigr]=-2i\partial_{x}\delta(x-y){\mathcal{T}}(y)$
$i\delta(\varphi-\theta){\mathcal{T}}^{\prime}(\theta)+\frac{c}{24\pi}i\delta^{\prime\prime\prime}(\varphi-\theta) +$$i\delta(x-y)\partial_{y}{\mathcal{T}}(y)+\frac{c}{24\pi}i\partial_{x}^{3}\delta(x-y)$
@@
Proj. rep. $\mathfrak{u}(\mathbf{X})$ $\mathfrak{u}(\mathbf{X})$ for $\mathbf{X}\in\operatorname{Vect}(S^{1})$ or $\operatorname{Vect}(S^{1})\otimes\mathbb{C}$ satisfying Eq. (A.2)
@@
$i\mathfrak{u}(\mathbf{X})=\int_{0}^{L}\textrm{d}x\,X(x){\mathcal{T}}(x)=\underset{m\in\mathbb{Z}}{\sum}\hat{X}_{m}\bigl(L_{-m}-\frac{c}{24}\delta_{m,0}\bigr)=\oint\textrm{d}\varphi\,X(\varphi){\mathcal{T}}(\varphi)$
$L_{n}=-\mathfrak{u}(\ell_{n})+\frac{c}{24}\delta_{n,0}$ for $\ell_{n}=-ie^{-in\varphi}\partial_{\varphi}$ satisfying $\ell_{m},\ell_{n}]=(m-n)\ell_{m+n}$
@@
Proj. rep. ${\cal U}(f)$ ${\cal U}(f)$ for $f\in\operatorname{Diff}(S^{1})$ satisfying Eq. (2.3)
@@
${\cal U}\bigl(e^{\lambda\mathbf{X}}\bigr)=e^{\lambda\mathfrak{u}(\mathbf{X})}$ for $\mathbf{X}\in\operatorname{Vect}(S^{1})$, $\lambda\in\mathbb{R}$
@@
Table 1: CFT formulas in angle coordinates and their counterparts in length coordinates.
@@
Some context and details for the algebra representation $\mathfrak{u}(\mathbf{X})$ in Table 1 are missing. To fill this gap, first note that Eq. (2.5) can be written as
@@
$\mathbf{X}=\sum_{m\in\mathbb{Z}}\hat{X}_{m}\,i\ell_{-m}\qquad\text{with}\qquad\ell_{m}\equiv-ie^{-im\varphi}\partial_{\varphi}$ (A.1)

### v2

@@ V.3 Return amplitudes and probabilities
A second application of Eq. (76) is to find the solution $U(t)|\psi(0)\rangle$ of the Schrödinger equation (46) for any initial condition $|\psi(0)\rangle$. As in Sec. IV, we choose $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$ to be an eigenstate of the initial inhomogeneous Hamiltonian, and investigate the return amplitude (44). This can again be done in terms of $2\times 2$ matrices, and the return amplitude can in fact be evaluated for any time $t$ (as opposed to $t=T$ at the end of one cycle). In terms of the convenient parameter $h_{k}\equiv h+\tfrac{c}{24}(k^{2}-1)$, one finds (see Appendix D)
$\langle\psi(0)|\psi(t)\rangle=e^{i\pi ck^{2}t/12L}u_{t}^{-2h_{k}/k}$ (83)
@@
$u_{t}$ $=\cos(k\Omega t/2)\cos(k\omega t/2)$ (84)
$\quad-2\pi\frac{\cosh(2\lambda)-L/T}{L\Omega}\sin(k\Omega t/2)\sin(k\omega t/2)$
$\quad+i\cosh(2\lambda)\biggl[\cos(k\Omega t/2)\sin(k\omega t/2)$
$\qquad+2\pi\frac{\operatorname{sech}(2\lambda)-L/T}{L\Omega}\sin(k\Omega t/2)\cos(k\omega t/2)\biggr]$
@@
expressed in terms of the frequency (78). The overall phase of $\langle\psi(0)|\psi(t)\rangle$ can be extracted by picking the branch of $\arg(u_{t})$ consistent with the limiting case $\lambda=0$, for which the phase is purely dynamical. This generalizes previous results for the return probability in $\mathfrak{sl}(2,\mathbb{R})$-driven CFT [99], which were restricted to piecewise-constant time-dependent Hamiltonians. Furthermore, the return probability was recently studied for continuous Floquet drives in [100], which relied on Floquet perturbation theory, while the result (83) is exact and uses no perturbation theory.
Note that the ground state ($h=0$) driven by deformations (71) with $k=1$ has $h_{1}=0$, owing to the definition above Eq. (83). As a result, the corresponding amplitude (83) is a pure phase. This is due to the fact that the ground state of a CFT is invariant under the ‘global conformal group’ of deformations (71) with $k=1$, as noted below Eqs. (30). (See also Footnote 11.) It is the reason why our simulations of $\mathrm{SL}(2,\mathbb{R})$ drives in Figs. 1(a) and 6 are carried out with $k=2$, as the latter is the lowest harmonic that gives rise to a nontrivial return probability for the ground state. [Incidentally, $\mathrm{SL}(2,\mathbb{R})$ drives with $k=1$ provide a good numerical check of the validity of CFT as an effective description of gapless lattice systems at low energies; see Appendix E.]
Adiabatic limit.
@@
As before, one is most interested in the return amplitude (44) after one period, in the adiabatic regime $L/T\ll 1$. Taking $t=T$ in Eq. (83) and expanding (84) in that regime yields
$\langle\psi(0)|\psi(T)\rangle=e^{i\Theta}\Bigl[1+O\bigl(L^{2}/T^{2}\bigr)\Bigr]$ (85)
@@
for $\Theta=2\pi\bigl(-h_{0}\frac{T}{L}+h_{k}\bigl[\cosh(2\lambda)-1\bigr]+h_{k}\frac{\sinh^{2}(2\lambda)}{2}\frac{L}{T}\bigr)+O(L^{2}/T^{2})$. The latter manifestly involves a dynamical phase, a Berry phase, a first nonadiabatic correction, and a rest term. The Berry phase, in particular, is consistent with Eq. (45) for a rotating drive in $\mathrm{SL}(2,\mathbb{R})$ [46].
@@
Note that the return amplitude (85) has unit norm up to $L^{2}/T^{2}$ corrections, so it is those corrections that are responsible for the nontrivial return probability $|\langle\psi(0)|\psi(T)\rangle|^{2}$. As in Sec. IV, the corrections are most conveniently found by expanding in powers of the adiabatic parameter (59) rather than $L/T$. One finds
$|\langle\psi(0)|\psi(T)\rangle|^{2}$
$=1-\delta^{2}\frac{2h_{k}}{k}\sinh^{2}(2\lambda)\sin^{2}(\pi k/\delta)+O\bigl(\delta^{3}\bigr),$ (86)
@@
which shows that there is an oscillating tail in the return probability beyond the adiabatic limit. This result is a special case of the general adiabatic expansion (63). A single Fourier mode $k$ now appears in the quantum metric (64), since in fact the vector field $\mathbf{Y}_{t}=Y_{t}(\varphi)\partial_{\varphi}$ of Eq. (62) is explicitly given by $Y_{t}(\varphi)=\sinh(2\lambda)\bigl[\sin(k\varphi)-\sin(k[\varphi+\Omega t])\bigr]$. The oscillations of Eq. (86) with $T$ can then be understood by referring to Fig. 5: the micromotion around the leading adiabatic approximation has a (high) frequency $\Omega$ of its own, which may or may not resonate with the (low) frequency $\omega$ of the drive. At resonance, when $\Omega T\in 2\pi\mathbb{Z}$, the initial and final rays coincide.
As at the end of Sec. IV, one should keep in mind that the driven Hamiltonian of a nonchiral CFT involves both right and left movers, as in Eq. (65). One should thus add to Eq. (72) the same expression with $\mathcal{T}$} replaced by $\overline{{\mathcal{T}}}$. This left-moving term only involves the generators $\{\bar{L}_{-k},\bar{L}_{0},\bar{L}_{k}\}$, which similarly generate an $\mathfrak{sl}(2,\mathbb{R})$ Lie algebra that commutes with the right-moving one, so that a matrix representation can again be used to solve the Schrödinger equation. In particular, the left-moving frequency $\overline{\Omega}$ in Eq. (69) becomes
$\overline{\Omega}=\frac{2\pi}{L}\sqrt{1+2\cosh(2\lambda)(L/T)+(L/T)^{2}},$ (87)
which is obtained from Eq. (78) by changing the sign in front of $\omega=2\pi/T$ wherever it appears. A key consequence is that the full nonchiral return probability (67)–(70) is the product of (86) with its left-moving counterpart:
$|\langle\psi(0)|\psi(T)\rangle|^{2}$ (88)
$\quad\begin{aligned} =1&-\delta^{2}\frac{2h_{k}}{k}\sinh^{2}(2\lambda)\sin^{2}(\pi k/\delta)\\ &-\bar{\delta}^{2}\frac{2\bar{h}_{k}}{k}\sinh^{2}(2\lambda)\sin^{2}(\pi k/\bar{\delta})+O\bigl(\delta^{3},\bar{\delta}^{3}\bigr),\end{aligned}$
where $\delta\equiv\omega/\Omega\neq\omega/\overline{\Omega}\equiv\bar{\delta}$ and $\bar{h}_{k}\equiv\bar{h}+\tfrac{\bar{c}}{24}(k^{2}-1)$. This is the nonchiral version of the chiral result (86). It is plotted (with a dashed red curve) in Fig. 1(a), and displays a characteristic oscillating convergence as $T/L\to\infty$. However, in contrast to the purely chiral case (86), it never quite reaches its asymptotic value due to the mismatch $\Omega\neq\overline{\Omega}$ [compare Eqs. (78) and (87)].
Comparison with lattice results.
@@
The comparison between CFT and a lattice model is displayed in Fig. 6. There, we plot return probabilities for a periodic free-fermion chain at half filling with $N$ sites and driven inhomogeneous hoppings
$J_{j}(t)=\cosh(2\lambda)-\sinh(2\lambda)\cos(k[2\pi j/N-\omega t]),$ (89)
@@
corresponding to the spin chain (2) for $\Delta=0$ with the same nearest-neighbor couplings. The return probability $|\langle\psi_{N}(0)|\psi_{N}(T)\rangle|^{2}$ of the lattice quantum state is computed numerically by approximating the evolution operator with a finite number of steps and using exact diagonalization for each time step; see Appendix E. The result exhibits remarkable agreement with the analytical formula obtained by taking the squared complex norm of Eq. (83), and multiplying it by itself with $\omega$ replaced by $\omega$ everywhere (including in $\Omega$). This is true even far away from the adiabatic regime, as long as the strength of the perturbation (controlled by $\lambda$) is sufficiently weak. Importantly, CFT results for the return probability are even applicable to lattice systems with a relatively small number of lattice sites, as shown in Fig. 6(b). This suggests that our dynamical probes for quantum geometry can realistically be implemented in quantum simulators with fewer than a hundred sites.
@@
Figure 6: (a) Return probability $|\langle\psi(0)|\psi(T)\rangle|^{2}$ after one period $T$ of the Hamiltonian (73), starting from the ground state $|\psi(0)\rangle$. The results are plotted as a function of $T/L$ for $\lambda=0.15,0.2,0.25$ (green to purple). (The parameters are $L=202$, $k=2$, $c=1$, $h=0$ and periodic boundary conditions are imposed.) We compare the exact CFT result with numerical lattice calculations for a free-fermion chain at half filling with $N=L$ sites (unit lattice spacing) and inhomogeneous hoppings of the form (89) (see Appendix E). The CFT result is the product of the squared complex norm of Eq. (83) with its left-moving counterpart; the latter is obtained from the same formula with $\omega$ replaced by $\omega$ everywhere—including in $\Omega$, which is thus replaced by $\overline{\Omega}$ in Eq. (87). (b) Finite-size scaling of the numerical lattice calculations close to the adiabatic limit, and comparison with the exact CFT result, for $k=2$, $\lambda=0.3$, and different system sizes $L$ ranging from 62 to 182 sites. As in Fig. 1, the agreement between numerics and analytical results is striking, now including the nonadiabatic region where $T/L\ll 1$.
VI Discussion
@@
This work explored the infinite-dimensional quantum geometry of 1+1D gapless many-body quantum systems through the lens of their low-energy dynamics, modeled as driven conformal field theories (CFTs). For slowly driven systems, we showed that the return probability probes the quantum metric through oscillations as a function of the driving frequency; see Figs. 1 and 6. In the complementary limit of arbitrarily fast but weak drives, absorption rates and linear-response coefficients respectively give access to the metric and the Berry curvature of the same quantum geometry.
These results are universal in the sense that they apply to any 1+1D gapless quantum system with an emergent conformal symmetry at low energies. They hold for both free and strongly interacting systems, and only depend on universal parameters such as the central charge of the effective theory. This is supported by a remarkable agreement between our field-theoretic results and numerics for gapless lattice models, opening a path to observe Virasoro quantum geometry in state-of-the-art quantum simulators.
Promising future research directions include:
@@
Geometric formulation of Floquet drives. It was recently understood that Floquet theory can be formulated from a quantum geometric perspective [25], by decomposing the Floquet evolution operator into a purely dynamical part (the ‘average-energy operator’) and a geometric part. It would be interesting to compute the average-energy operator for our periodic drives, and understand the link between the Virasoro Berry phases (45) [46] and those of [25].
@@
Higher dimensions. Our study of the quantum geometry of 1+1D gapless many-body quantum systems relied on the Virasoro algebra, allowing us to probe an infinite-dimensional parameter space (18). While this algebraic structure is absent in higher-dimensional CFTs, one may still design nonequilibrium probes of the quantum geometry associated with the global conformal group SO(d+1,2)(d+1,2) of gapless many-body systems in dd+1 spacetime dimensions. Indeed, given that the approach of this work holds for any continuous symmetry group, it could naturally be applied to higher-dimensional CFTs.
@@
Quantum geometry of mixed states. Our work concerned nonequilibrium probes of the quantum geometry for pure states, specifically Virasoro coherent states (23). The relevant quantum metric then was (the pullback of) the Fubini-Study metric. A natural next step is to consider mixed states, and dynamically probe their quantum geometry. In the strict adiabatic regime, one may expect a suitable generalization of our protocol to feature the Uhlmann phase [101, 102], as a mixed-state counterpart of the Berry phase. Crucially though, the corresponding metric is not unique [103], so subleading corrections in the adiabatic regime would probe different metrics depending on the choice of distance measure. Examples include the Bures metric and the Kubo-Mori metric; the former was recently considered in [104, 105], and the latter was studied in [42] for thermal states in driven CFT.
Acknowledgments.
@@
We are especially grateful to Bruno Mera for useful comments on the first version of this paper. We also thank Mathieu Beauvillain, Marin Bukov, Adolfo del Campo, Raphaël Ducatez, Benoit Estienne, Gian Michele Graf, Gregor Jotzu, Shinsei Ryu, and Erik Tonni for inspiring discussions. B.L. acknowledges financial support from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 214461). P.M. acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002).
@@ Appendix A Brief dictionary of CFT formulas
Virasoro algebra ${\mathcal{T}}_{\pm}(\varphi)=\frac{1}{2\pi}\underset{n\in\mathbb{Z}}{\sum}e^{\pm in\varphi}\bigl(L^{\pm}_{n}-\frac{c}{24}\delta_{n,0}\bigr)$ ${\mathcal{T}}_{\pm}(x)=\bigl(\frac{2\pi}{L}\bigr)^{2}{\mathcal{T}}_{\pm}(\varphi)=\frac{2\pi}{L^{2}}\underset{n\in\mathbb{Z}}{\sum}e^{\pm i\frac{2\pi}{L}nx}\bigl(L^{\pm}_{n}-\frac{c}{24}\delta_{n,0}\bigr)$
$L^{\pm}_{m}=\oint\textrm{d}\varphi\,e^{\mp im\varphi}\,{\mathcal{T}}_{\pm}(\varphi)+\frac{c}{24}\delta_{m,0}$ satisfying Eq. (15)
@@
$\bigl[{\mathcal{T}}_{\pm}(\varphi),{\mathcal{T}}_{\pm}(\theta)\bigr]=\mp 2i\delta^{\prime}(\varphi-\theta){\mathcal{T}}_{\pm}(\theta)$ $\bigl[{\mathcal{T}}_{\pm}(x),{\mathcal{T}}_{\pm}(y)\bigr]=\mp 2i\partial_{x}\delta(x-y){\mathcal{T}}_{\pm}(y)$
$\pm i\delta(\varphi-\theta){\mathcal{T}}_{\pm}^{\prime}(\theta)\pm\frac{c}{24\pi}i\delta^{\prime\prime\prime}(\varphi-\theta)$ $\pm i\delta(x-y)\partial_{y}{\mathcal{T}}_{\pm}(y)\pm\frac{c}{24\pi}i\partial_{x}^{3}\delta(x-y)$
@@
Proj. anti- Hermitian rep. $\mathfrak{u}_{\pm}(\mathbf{X})$ $\mathfrak{u}_{\pm}(\mathbf{X})$ for $\mathbf{X}\in\operatorname{Vect}(S^{1})$ or $\operatorname{Vect}(S^{1})\otimes\mathbb{C}$ satisfying Eq. (91)
@@
$\mathfrak{u}_{\pm}(\mathbf{X})=\partial_{\epsilon}\,{\cal U}_{\pm}(\mathbb{I}+\epsilon\mathbf{X})\big|_{\epsilon=0}=\mp i\oint\textrm{d}\varphi\,X(\varphi){\mathcal{T}}_{\pm}(\varphi)=\mp i\underset{m\in\mathbb{Z}}{\sum}\hat{X}_{m}\bigl(L^{\pm}_{\mp m}-\frac{c}{24}\delta_{m,0}\bigr)=\mp i\int_{0}^{L}\textrm{d}x\,X(x){\mathcal{T}}_{\pm}(x)$
$L^{\pm}_{n}=-\mathfrak{u}_{\pm}(\ell_{\mp n})+\frac{c}{24}\delta_{n,0}$ for $\ell_{n}=-ie^{-in\varphi}\partial_{\varphi}$ satisfying $\ell_{m},\ell_{n}]=(m-n)\ell_{m+n}$
@@
Proj. unitary rep. ${\cal U}_{\pm}(f)$ ${\cal U}_{\pm}(f)$ for $f\in\operatorname{Diff}(S^{1})$ satisfying ${\cal U}_{\pm}(f)\,{\cal U}_{\pm}(g)=e^{\pm ic\hskip 0.7pt\mathsf{C}(f,g)}\,{\cal U}_{\pm}(f\circ g)$ with $\mathsf{C}(f,g)$ in Eq. (10)
@@
${\cal U}_{\pm}\bigl(e^{\epsilon\mathbf{X}}\bigr)=e^{\epsilon\mathfrak{u}_{\pm}(\mathbf{X})}$ for $\mathbf{X}\in\operatorname{Vect}(S^{1})$, $\epsilon\in\mathbb{R}$ $\implies$ ${\cal U}_{\pm}(\mathbb{I}+\epsilon\mathbf{X})=1+\epsilon\mathfrak{u}_{\pm}(\mathbf{X})+o(\epsilon)$
@@
Table 1: Useful formulas for nonchiral 1+1D CFT in angle coordinates and their counterparts in length coordinates.
This appendix complements the concise review of CFT and Virasoro symmetry in Sec. II, extended to the nonchiral case of Eq. (65). For convenience, we change notation to subscripted ++ or - to respectively indicate right or left movers. This includes e.g. ${\mathcal{T}}_{+}(\varphi^{-})={\mathcal{T}}(\varphi^{-})$ and ${\mathcal{T}}_{-}(\varphi^{+})=\overline{{\mathcal{T}}}(\varphi^{+})$, each depending only on its respective light-cone coordinate $\varphi^{\pm}$ defined above Eq. (7). However, we set $c=\bar{c}$ for simplicity.
Table 1 summarizes useful formulas for nonchiral 1+1D CFT in angle coordinates $\varphi$, $\theta$ vs. length coordinates $x\equiv\frac{L}{2\pi}\varphi$, $y\equiv\frac{L}{2\pi}\theta$. Quantities are distinguished by their arguments, $\delta(\varphi)$ and $\delta(x)$ respectively denote the $2\pi$- and $L$-periodic delta functions, a prime denotes differentiation with respect to $\varphi$, and $\oint\textrm{d}\phi$ is shorthand for $\int_{0}^{2\pi}\textrm{d}\phi$. Moreover, $\operatorname{Diff}(S^{1})$ denotes the universal cover of the group of orientation-preserving diffeomorphisms of the circle $S^{1}$. $\operatorname{Vect}(S^{1})$ denotes the corresponding Lie algebra of real, smooth vector fields on $S^{1}$, and $\operatorname{Vect}(S^{1})\otimes\mathbb{C}$ denotes its complexification.
@@
Some context and details for the algebra representation $\mathfrak{u}_{\pm}(\mathbf{X})$ in Table 1 are missing. To fill this gap, first note that Eq. (11) can be written as
@@
$\mathbf{X}=\sum_{m\in\mathbb{Z}}\hat{X}_{m}\,i\ell_{-m}\quad\text{with}\quad\ell_{m}\equiv-ie^{-im\varphi}\partial_{\varphi}$ (90)